What is the equation of the line passing through (-3, 8) that is perpendicular to the line y = 7?
thanks so much

Answer by cira554(7)

equation of a line is usually in the form y=mx+c where m is the slope
in our equation y=7,the 7 is our slope.
for perpendicular lines the product of their slopes is -1
that is m1*m2=-1
our m1 is 7 so our m2 will be -1/7
assuming our line passes through the point(x,y) and(-3,8) for us to get the slope we are supposed to get change in y/change in x
hence y-8/x-(-3)=-1/7
crossmultiply and get 7y-56=-x-3
y=-x/7+53/7 which is the equation to the line